OpenStudy (anonymous):

find the x-intercepts of the parabola with vertex (6,-5) and y-intercept (0,175).

OpenStudy (anonymous):

can anyone help quick?

OpenStudy (anonymous):

@djones3507 To find the x value of the vertex, use -b/2a, so our first equation is: -b/2a = 6 To find the y value of the vertex, substitute the x value from -b/2a into the quadratic formula. Remember, we already know c the c value (the y intercept) with is 175, so our second equation is a(6²) + b(6) + 175 = -5 Solve for b in the second equation in terms of a: b = -6a - 30 Substitute into our first equation: 6a + 30/2a = 6 6a + 30 = 12a 6a = 30 a = 5 Now, we use the second equation to find our b value b = -6(5) - 30 b = -60 This means our quadratic is : 5x² - 60x + 175 = 0 This simplifies to x² - 12x + 35 = 0 Factor as: (x - 7)(x - 5) = 0 The x intercepts are (5,0) and (7,0)

OpenStudy (anonymous):

:))